OpenStudy (anonymous):

How would you solve/approach this: sin−1(sin(7π/3))

OpenStudy (anonymous):

Sin-1(sin x) =x

myininaya (myininaya):

yes true!

OpenStudy (anonymous):

now i have sin−1(sin(7π/3)=7π/3 do i divide each side by sinx?

myininaya (myininaya):

remember though we want x to be between -pi/2 and pi/2

myininaya (myininaya):

recall sin(pi/3)=sin(7pi/3) so sin^{-1}(sin(7pi/3))=sin^{-1}(sin(pi/3))=pi/3

myininaya (myininaya):

oh it is sin-1(sin(7pi/3))?

OpenStudy (anonymous):

how do you know pi/3 = 7pi/3?

myininaya (myininaya):

this doesn't make any sense then

myininaya (myininaya):

sin(pi/3)=sin(7pi/3)

myininaya (myininaya):

i didn't say pi/3=7pi/3

OpenStudy (anonymous):

your job is to find the number between \[-\frac{\pi}{2},\frac{\pi}{2}\] whose sine is given. in other words look on the right side of the unit circle, and find an agle coterminal with \[\frac{7\pi}{3}\]

OpenStudy (anonymous):

*angle

myininaya (myininaya):

right since the range of sine inverse is (-pi/2,pi/2)

OpenStudy (anonymous):

to be pedantic is it \[[-\frac{\pi}{2}, \frac{\pi}{2}]\]

OpenStudy (anonymous):

is pi/6 coterminal to 7pi/3?

myininaya (myininaya):

lol very good satellite

myininaya (myininaya):

what is 7pi/3-2pi?

OpenStudy (anonymous):

no because your denominator is wrong. should be \[\frac{7\pi}{3}-2\pi=\frac{\pi}{3}\]

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

oh! i see

OpenStudy (anonymous):

math really isn't my subject :/ but so once i have sin-1(sin(pi/3)) then since that equals x, does it evaluate to pi/3?

OpenStudy (anonymous):

you could also do this problem the donkey way, but just saying \[\sin(\frac{7\pi}{3})=\frac{\sqrt{3}}{2}\] and then finding \[\sin^{-1}(\frac{\sqrt{3}}{2})\]which is \[\frac{\pi}{3}\] but that is the silly way

OpenStudy (anonymous):

yes, answer is \[\frac{\pi}{3}\]

OpenStudy (anonymous):

why is that the silly way just wondering? lol

OpenStudy (anonymous):

what I don't understand is how you find what sin(7pi/3) equals?

OpenStudy (anonymous):

how do i know it?

OpenStudy (anonymous):

like how do you solve that?

OpenStudy (anonymous):

i find \[\frac{7\pi}{3}\] on the unit circle and look at the second coordinate that is the sine first coordinate is cosine

OpenStudy (anonymous):

look at the unit circle on last page of cheat sheet. find the angle you are looking for and then look at the second coordinate

OpenStudy (anonymous):

isn't that -1/2

OpenStudy (anonymous):

first coordinate is 1/2

OpenStudy (anonymous):

because isn't the coordinate is (-sqrt3/2, -1/2)?

OpenStudy (anonymous):

by counting out around the circle you will see that \[\frac{7\pi}{3}\] is coterminal with \[\frac{\pi}{3}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

this is what confuses me. 7pi/3 isn't on the unit circle.

OpenStudy (anonymous):